A semi-elliptical archway over a one-way road has a height of 9 ft and width of 36 ft. Your van has a width of 12 ft and height of 8 ft. Will your van be able to clear the opening of the archway?

1
Expert's answer
2021-11-25T13:59:05-0500

Suppose that the eqaution of of our semi-ellipse is


"\dfrac{x^2}{18^2}+\dfrac{y^2}{9^2}=1, y\geq 0"

If "x=6"


"\dfrac{6^2}{18^2}+\dfrac{y^2}{9^2}=1"

"y^2=81-9, y\geq0"

"y=6\sqrt{2}"

Since "6\sqrt{2}>8," then van will be able to clear the opening of the archway.